# Evaluating and improving wave and non-wave stress parametrisations for oceanic flows

### v1

@@ 2.2 Case 2: Steady Flow
We now consider a constant steady flow $U$ in the $x$-direction. Under such a flow, the internal waves formed are called lee waves. These lee waves have zero Eulerian frequency (i.e. are steady at a fixed point), and are formed with intrinsic frequency depending on the topography and the magnitude of $U$. This leads to some added complexities compared to the tidal case, where the internal waves have the same intrinsic frequency as the tide. In particular, for steady flows there is a greater dependency on the Froude number
@@
As in the tidal case, we also focus on two parametrisations for steady flows, treating the cases of small and large Froude numbers separately. The first, which is based on the linear theory of Bell (1975), is an idealised theory assuming lee waves form without the effects of a high Froude number. The second, is a semi-empirical parametrisation due to Klymak et al. (2010, hereafter, KLP2010). The KLP2010 parametrisation focuses instead on a regime where most of the flow is blocked, and unable to surmount the topography to form lee waves. These two parametrisations therefore look at different aspects of the wave and non-wave stress, and so both are useful in obtaining a wholistic parametrisation.
@@ 2.2.1 Bell parametrisation for steady flows
In the work of Bell (1975), an expression for the time-mean stress due to lee waves is provided. In three dimensions, their expression for the ($x$-component of) stress is
@@
where $\alpha$ is as in (6) and the integrals are taken over all real values of $k$ and $\ell$ under the restriction $|f|\leq|kU|<\infty$ for hydrostatic ($\alpha=0$) and $|f|\leq|kU|\leq N$ for non-hydrostatic ($\alpha=1$). As we are taking the flow speed to be constant, the lee wave stress is non-zero in the time-mean and there is not an equivalent bottom-trapped spring force as in the SAH2020 tidal parametrisation theory. However, it is worth noting that (2.2.1) can readily be obtained by repeating the arguments in SAH2020 for a steady flow. The two-dimensional equivalent of (2.2.1) is then
@@
whenever $Fr\geq Fr_{c}$. The standard values of $Fr_{c}$ used are 0.7 in two dimensions, and 0.4 in three dimensions (Nikurashin et al., 2014).
@@ 2.2.2 Klymak, Legg and Pinkel parametrisation
In KLP2010, a form drag parametrisation is obtained for when the Froude number $Fr$ is very large, whereby blocking effects are the dominant form of stress. Their parametrisation is derived using numerical simulations for two-dimensional Gaussian hills under the hydrostatic assumption and $f=0$. The full expression for the form drag in KLP2010 is
@@
For $Fr\gg 1$, the second and third terms in (18) are small, so that this parametrisation is nearly equivalent to the linear lee wave stress given in (16). The main difference is the scaling of $U$ by $H/(H-h)$, which represents the increase in velocity as the flow moves over the narrower channel above the hill. This suggests that when $f=0$, an approximate $Nh_{0}^{2}U$ scaling could be accurate for both low and high Froude numbers.
@@ 2.3 Case 3: Mixed Flow
where $\phi$ denotes a phase shift. A naive parametrisation would be to assume that the mean (steady) and tidal components of the flow do not interact. That is, $F_{m}$ is independent of $U_{\text{tidal}}$ and $F_{p}$ is independent of $U_{\text{mean}}$, so that $F_{m}$ and $F_{p}$ can be directly computed from the parametrisations in the steady and tidal cases respectively. In terms of a more sophisticated parametrisation, we now discuss a more general version of the Bell (1975) theory.
@@
Although we previously discussed the Bell (1975) theory for steady flows, Bell also gives a more general expression for mixed flows. Namely, in three dimensions, Bell gives the following formula for the $x$-component of the time-mean stress (cf. (2.2.1))
@@
×$\qquad\qquad\qquad\qquad\times J_{n}^{2}\left(\frac{kU_{\text{tidal}}}{\omega}\right)\mathrm{d}k\mathrm{d}\ell,$ (22)
@@
×$\qquad\qquad\qquad\qquad\times J_{n}^{2}\left(\frac{kU_{\text{tidal}}}{\omega}\right)\mathrm{d}k.$ (23)
There are some notable features of (2.3). Firstly, if $U_{\text{tidal}}=0$ then (2.3) reduces to the same expression in the steady flow case (2.2.1). Moreover, if $U_{\text{mean}}=0$, we have $F_{m,3\text{d}}=F_{m,2\text{d}}=0$. We also note that $J_{0}(0)=1$ and $J_{n}(0)=0$ for $n>0$ but in general $|J_{0}(x)|\leq 1$ and $|J_{n}(x)|\geq 0$. This can be interpreted as the lee wave energy ($n=0$) being transferred to higher modes ($n>0$) in the presence of a tidal flow (Shakespeare, 2020). However, since $|J_{n}(x)|\ll|J_{0}(x)|$ for $n\geq 1$ and small $x$, the time-mean stress predicted by (2.3) and (2.3) is often not too different from the case of a purely steady flow.
Finally we note that the Bell parametrisation (2.3) only gives a small glimpse of understanding on how to parametrise stress for mixed flows. Like the Bell parametrisation for a steady flow, (2.3) only applies in the linear regime, where the topographical height is assumed to be small. Bell’s parametrisation also does not tell us anything about the periodic component of stress $F_{p}$. Thus, we will primarily analyse our simulation data for $F_{p}$ by simply comparing it with the tidal case where $U_{\text{mean}}=0$. However, one should expect some impact of the mean flow on the periodic stress given that the mean flow will act to Doppler shift the tidal frequency $\omega$.
@@ 3 Numerical Model
We use Oceananigans (Ramadhan et al., 2020; Wagner et al., 2025) to model flow over an isolated Gaussian hill in two or three dimensions. Importantly, the GPU capability of Oceananigans allows us to run several hundred simulations whilst using only a modest amount of computational resources (see also Silvestri et al. (2024)). In what follows, we focus on the setup in three dimensions, as the two-dimensional simulations are essentially the same but with only one grid cell in the $y$-direction and improvements in the resolution. For more precise details of the setup see Johnston (2025), which contains the base code used for each simulation.
@@
By default, our domain was set to be $L_{x}=L_{y}=500\ \text{km}$ long and wide, $H=1.5\ \text{km}$ deep and centred on a symmetric Gaussian hill $h(x,y)=h_{0}e^{-(x^{2}+y^{2})/(2W^{2})}$ with height $h_{0}$ and half-width $W=5\ \text{km}$. Our domain was periodic in the horizontal $x$-$y$ directions, and bounded in the vertical. To capture the features near the hill, the domain grid was stretched to give a horizontal spacing of $\Delta x=\Delta y\approx 700\ \text{m}$ around the hill, smoothly varied to $\Delta x=\Delta y\approx 2100\ \text{m}$ at the edge of the domain. Similarly, the vertical grid spacing was set to be $\Delta z\approx 7.5\ \text{m}$ in the bottom half of the domain, smoothly varied to $\Delta z\approx 22.5\ \text{m}$ near the surface.
To keep the hill isolated, we artificially damped wave features at the horizontal boundaries by introducing a “sponge layer” in the horizontal regions $[-L_{x}/2,-L_{x}/4]$, $[L_{x}/4,L_{x}/2]$, $[-L_{y}/2,-L_{y}/4]$ and $[L_{y}/4,L_{y}/2]$. In particular, the horizontal diffusivity and horizontal viscosity coefficients were smoothly varied from 0 near the hill, to $2000\ \text{m}^{2}\text{s}^{-1}$ in the outer regions of the domain. Note that everywhere the above-mentioned coefficients were the same. Similarly, the buoyancy and horizontal velocity were relaxed to their initial conditions at the horizontal boundaries to prevent wave features from propagating across the periodic boundary and obfuscating the results.
@@
For each simulation, we specified constant values for background stratification ($N$), hill height ($h_{0}$), hill width ($W$), ocean depth ($H$), Coriolis frequency ($f$), and tidal frequency ($\omega$) if required. Depending on the whether we were concerned with a tidal, mean or mixed case, we also specified the flow speeds $U_{\text{mean}}$ and $U_{\text{tidal}}$ (see (20)). The velocity is initialised as $\bm{u}(t=0)=(U_{\text{mean}}+U_{\text{tidal}},0,0)$ and the buoyancy is initialised with $b(t=0)=N^{2}z$. The velocity is then forced analogously to the simulations by Shakespeare et al. (2021) or Klymak (2018) to maintain a background zonal velocity $U(t)=U_{\text{mean}}+U_{\text{tidal}}\cos(\omega t)$. In particular, a forcing of $F_{y}=fU_{\text{mean}}$ is added to the $y$-momentum equations and a forcing of
$F_{x}=\frac{f^{2}-\omega^{2}}{\omega}U_{\text{tidal}}\sin(\omega t)$ (24)
@@
is added to the $x$-momentum equations. Our general procedure was to run a test using reference values for $N,\ h_{0},\ W,\ H,\ f,\ \omega,\ U_{\text{mean}}$ and $U_{\text{tidal}}$, before varying each parameter individually to determine their impact on the stress. For the purely tidal or steady flow tests, these default parameters are listed in Table 1. Note that the bottom-trapped tests are run at $90^{\circ}S$ which, although unphysical, is fine in a simulation and allows us to study a regime whereby bottom-trapped effects truly dominate. For the mixed flow case, the parameters will either be the same as the tidal case with a mean velocity of $U_{\text{mean}}=0.1\ $\mathrm{m}\,\mathrm{s}^{-1}$$ added, or the same as the mean case with a tidal velocity of $0.1\cos(\omega t)\ $\mathrm{m}\,\mathrm{s}^{-1}$$ added. This will be made clear in the text. Also note that the default values of $h_{0}$ and $H$ vary slightly between the two and three-dimensional tests. This is due to it being more difficult to achieve sufficient resolution for small hills in the three-dimensional case.
The $x$-component of the stress was directly computed as in  (2):
@@
where $p_{h}(x,y)$ is the pressure corresponding to the grid cell directly above the hill at horizontal position $(x,y)$.
Table 1: The default parameter values used for the tests with a purely tidal or steady flow in either two dimensions (2d) or three dimensions (3d). For the tidal flow tests, the choice of $\omega$ and $f$ (and thus latitude) will be made clear in the text. The parameters for the steady flow tests vary depending on whether we are interested in a low or high Froude number $Fr=Nh_{0}/U_{\text{mean}}$ regime.
Tidal Steady $(Fr\lesssim 1)$ Steady ($Fr\gg 1$)
$N$ $(\mathrm{s}^{-1})$ $0.004$ $0.002$ $0.005$
$h_{0}$ $(\mathrm{m})$ $100$ $20$ (2d) or $30$ (3d) $500$
$W$ $(\mathrm{m})$ $5000$ $5000$ $5000$
$H$ $(\mathrm{m})$ $1500$ $1500$ (2d) or $1250$ (3d) $1500$
$f$ $(\mathrm{s}^{-1})$ $-5\cdot 10^{-5}$ or $-1.46\cdot 10^{-4}$ $-2.53\cdot 10^{-5}$ 0
Latitude $(^{\circ}S)$ 20 or 90 10 0
$\omega$ $(\mathrm{s}^{-1})$ $1.4\cdot 10^{-4}$ or $7.3\cdot 10^{-5}$- --
$U_{\text{mean}}$ ($\mathrm{m}\,\mathrm{s}^{-1}$)$ 00 $0.2$ $0.1$
$Fr=Nh_{0}/U_{\text{mean}}$- $0.2$ (2d) or $0.3$ (3d) $25$
$U_{\text{tidal}}$ ($\mathrm{m}\,\mathrm{s}^{-1}$)$ $0.1$ 00 00
For efficiency, all simulations were run in hydrostatic mode. The unimportance of non-hydrostatic processes for the stress is reflected by the parametrisations in Section 2, whereby the hydrostatic ($\alpha=0$) and non-hydrostatic ($\alpha=1$) parametrisations are essentially the same provided the buoyancy frequency $N$ is large compared to the wave frequency. Each three-dimensional tidal test was run for 200 hours, whereas each three-dimensional steady test was run for 100 hours. We found that 100 hours was generally long enough to reach a near-equilibrium state, with the slightly longer test time given for tidal tests as to have more full tidal cycles to perform a statistical analysis on. Since two-dimensional tests were much less computationally expensive, we decided to run them for even longer (750 hours) to improve accuracy and detect any notable effects that only appeared at later times. All simulations used time-step of 7.5 minutes, and an output interval of 15 minutes.

### v3

@@ 2.2 Case 2: Steady Flow
We now consider a constant steady flow $U$ in the $x$-direction. Under such a flow, the internal waves formed are called lee waves. These lee waves have zero Eulerian frequency (i.e., are steady at a fixed point), and are formed with intrinsic frequency depending on the topography and the magnitude of $U$. This leads to some added complexities compared to the tidal case, where the internal waves have the same intrinsic frequency as the tide. In particular, for steady flows there is a greater dependency on the Froude number
@@
As in the tidal case, we also focus on two parametrisations for steady flows, treating the cases of small and large Froude numbers separately. The first, which is based on the linear theory of Bell (1975), is an idealised theory assuming lee waves form without the effects of a high Froude number. The second, is a semi-empirical parametrisation due to Klymak et al. (2010, hereafter, KLP2010). The KLP2010 parametrisation focuses instead on a regime where most of the flow is blocked, and unable to surmount the topography to form lee waves. These two parametrisations therefore look at different aspects of the wave and non-wave stress, and so both are useful in obtaining a wholistic parametrisation. Of course, there is also the transitional case where Froude number is of a moderate size, say $Fr\approx 1-10$. To limit the scope of our analysis we do not explore this case in depth here, but refer the reader to recent attempts at developing a hybrid parametrisation (Perfect et al., 2020; Klymak et al., 2021; Baker and Mashayek, 2022).
@@ 2.2.1 Bell parametrisation for steady flows
In the work of Bell (1975), an expression for the time-mean stress due to lee waves is provided. In three dimensions, his expression for the $x$-component of stress is
@@
where $\alpha$ is as in (5) and the integrals are taken over all real values of $k$ and $\ell$ under the restriction $|f|\leq|kU|<\infty$} for hydrostatic ($\alpha=0$) and $|f|\leq|kU|\leq N$ for non-hydrostatic ($\alpha=1$) conditions. As we are taking the flow speed to be constant, the lee wave stress is non-zero in the time-mean and there is not an equivalent bottom-trapped spring force as in the SAH2020 tidal parametrisation theory. However, it is worth noting that (2.2.1) can readily be obtained by repeating the arguments in SAH2020 for a steady flow. The two-dimensional equivalent of (2.2.1) is then
@@
whenever $Fr\geq Fr_{c}$. The standard values of $Fr_{c}$ used are 0.7 in two dimensions, and 0.4 in three dimensions. These values were derived empirically from experimental observations (Aguilar and Sutherland, 2006) and simulation data (Nikurashin et al., 2014).
@@ 2.2.2 Klymak, Legg and Pinkel parametrisation
In KLP2010, a form stress parametrisation is obtained for when the Froude number $Fr$ is very large, whereby blocking effects are the dominant form of stress. Their parametrisation is derived using numerical simulations for two-dimensional Gaussian hills under the hydrostatic assumption and $f=0$. The full expression for the form stress in KLP2010 is
@@
For $Fr\gg 1$, the second and third terms in (18) are small, so that this parametrisation is nearly equivalent to the linear lee wave stress given in (16). The main difference is the scaling of $U$ by H/(Hh)H/(H-h), which represents the increase in velocity as the flow moves over the narrower channel above the hill. This suggests that when $f=0$, an approximate $Nh_{0}^{2}U$ scaling could be accurate for both low and high Froude numbers.
@@ 2.3 Case 3: Mixed Flow
where $\phi$ denotes a phase shift. A naive parametrisation would be to assume that the mean (steady) and tidal components of the flow do not interact (although we know this is not the case; e.g., Shakespeare, 2020). Nonetheless, if we assume that $F_{m}$ is independent of $U_{\text{tidal}}$ and $F_{p}$ is independent of $U_{\text{mean}}$, then $F_{m}$ and $F_{p}$ can be directly computed from the parametrisations in the steady and tidal cases respectively. There does not appear to be any well-developed parametrisation for $F_{p}$ in the literature (Bell, 1975; Shakespeare, 2020, only consider the energy flux), so we primarily compare our simulation data to this naive parametrisation for $F_{p}$. Certainly though, one should expect some impact of the mean flow on the periodic stress given that the mean flow will act to Doppler shift the tidal frequency. For the mean component $F_{m}$ however, there is some existing theory available, which we now discuss.
@@
Although we previously discussed the Bell (1975) theory for steady flows, Bell also gives a more general expression for the mean stress in mixed flows. Shakespeare (2020) extended this expression to include multiple tidal frequencies, but we will only consider the simpler Bell formulation here. In three dimensions, Bell gives the following formula for the $x$-component of the time-mean stress (cf. (2.2.1))
@@
$\qquad\qquad\qquad\qquad\times J_{n}^{2}\left(\frac{kU_{\text{tidal}}}{\omega}\right)\mathrm{d}k\mathrm{d}\ell,$ (22)
@@
$\qquad\qquad\qquad\qquad\times J_{n}^{2}\left(\frac{kU_{\text{tidal}}}{\omega}\right)\mathrm{d}k.$ (23)
There are some notable features of (2.3). Firstly, if $U_{\text{tidal}}=0$ then (2.3) reduces to the same expression in the steady flow case (2.2.1). Moreover, if $U_{\text{mean}}=0$, we have $F_{m,3\text{d}}=F_{m,2\text{d}}=0$. We also note that, provided $x$ is not too large, $J_{0}(x)$ decreases from $1$ whilst $J_{n}(x)$ increases from 0 for $n\geq 1$. This can be interpreted as the lee wave energy ($n=0$) being transferred to higher modes ($n\geq 1$) in the presence of a tidal flow (Shakespeare, 2020). However, in most cases $J_{n}(x)$ ($n\geq 1$) is sufficiently small so that the time-mean stress predicted by (2.3) and (2.3) is not too different from the case of a purely steady flow.
Like the Bell parametrisation for a steady flow, (2.3) and (2.3) are derived in the linear regime, where the Froude number is assumed to be small. Thus, for large Froude numbers, where this theory is invalid, we will primarily compare our simulation data to the naive parametrisation which assumes that the tidal and mean components of the flow do not interact.
@@ 3 Numerical Model
We use the Julia library Oceananigans (Ramadhan et al., 2020; Wagner et al., 2025) to model flow over an isolated Gaussian hill in two or three dimensions. Importantly, the GPU-capability of Oceananigans allows us to run several hundred simulations whilst using only a modest amount of computational resources (see also Silvestri et al. (2025)). In what follows, we focus on the setup in three dimensions, as the two-dimensional simulations are essentially the same but with only one grid cell in the $y$-direction and improvements in the resolution. For more precise details of the setup see Johnston et al. (2025), which contains the base code used for each simulation.
@@
By default, our domain was set to be $L_{x}=L_{y}=500\ \text{km}$ long and wide, H=1.5kmH=1.5\ $\text${km} deep and centred on a symmetric Gaussian hill $h(x,y)=h_{0}\,e^{-(x^{2}+y^{2})/(2W^{2})}$ with height $h_{0}$ and width W=5kmW=5\ $\text${km}. Our domain was periodic in the horizontal $x$-$y$ directions, and bounded in the vertical. To capture the features near the hill, the domain grid was stretched to give a horizontal spacing of $\Delta x=\Delta y\approx 700\ \text{m}$ around the hill, smoothly varied to $\Delta x=\Delta y\approx 2100\ \text{m}$ at the edge of the domain. Similarly, the vertical grid spacing was set to be $\Delta z\approx 7.5\ \text{m}$ in the bottom half of the domain, smoothly varied to $\Delta z\approx 22.5\ \text{m}$ near the surface. The uneven bathymetry is implemented in Oceananigans as an immersed boundary.
To keep the hill isolated, we artificially damped wave features at the horizontal boundaries by introducing a “sponge layer” in the horizontal regions $L_{x}/2,-L_{x}/4]$, $L_{x}/4,L_{x}/2]$, $L_{y}/2,-L_{y}/4]$ and $L_{y}/4,L_{y}/2]$. The sponge layer, in $x$, is:
$\mathrm{sponge}(x)=\frac{1}{2}\mathrm{Erf}\left(\frac{x-L_{x}/4}{L_{x}/16}\right)+\frac{1}{2}\mathrm{Erf}\left(\frac{-x-L_{x}/4}{L_{x}/16}\right)+1,$ (24)
[…]
@@
For each simulation, we specified constant values for background stratification ($N$), hill height ($h_{0}$), hill width ($W$), ocean depth ($H$), Coriolis frequency ($f$), and tidal frequency ($\omega$) if required. Depending on whether we were concerned with a tidal, mean or mixed case, we also specified the flow speeds $U_{\text{mean}}$ and $U_{\text{tidal}}$ (see (20)). The velocity is initialised as
$\bm{u}(t=0)=(U_{\text{mean}}+U_{\text{tidal}})\hat{\bm{x}}$, (25)
and buoyancy is initialised with a linear constant stratification:
$b(t=0)=N^{2}z.$ (26)
The horizontal momentum equations are forced analogously to the simulations by Klymak (2018) and Shakespeare et al. (2021a) to maintain a background zonal velocity $U(t)=U_{\text{mean}}+U_{\text{tidal}}\cos(\omega t)$}. In particular, the $x$-momentum equation is forced with:
$\frac{f^{2}-\omega^{2}}{\omega}U_{\text{tidal}}\sin(\omega t),$ (27)
and the $y$-momentum equation with $fU_{\text{mean}}$.
@@
Our general procedure was to run a test using reference values for $N,\ h_{0},\ W,\ H,\ f,\ \omega,\ U_{\text{mean}}$ and $U_{\text{tidal}}$, before varying each parameter individually to determine their impact on the stress. For the purely tidal or steady flow tests, these default parameters are listed in Table 1. Note that the bottom-trapped tests are run at $90^{\circ}S$ which, although unphysical, is fine in a simulation and allows us to study a regime whereby bottom-trapped effects truly dominate. For the mixed flow case, the parameters will either be the same as the tidal case with a mean velocity of $U_{\text{mean}}=0.1\ $\mathrm{m}\,\mathrm{s}^{-1}$\$ added, or the same as the mean case with a tidal velocity of $0.1\cos(\omega t)\ $\mathrm{m}\,\mathrm{s}^{-1}$\$ added. This will be made clear in the text. Also note that the default values of $h_{0}$ and $H$ vary slightly between the two- and three-dimensional tests. This is due to it being more difficult to achieve high resolution for small hills in the three-dimensional case.
In each test, the main quantity we computed was the $x$-component of the stress, given as in (2) by:
@@
where $p_{h}(x,y)$ is the bottom pressure, i.e., the pressure values that corresponds to the grid cell directly above the hill at horizontal position $(x,y)$.
The default values in columns 1 and 2 of Table 1 were chosen to be within the usual linear regimes. This allowed us to clearly see any non-linear effects that arose when the parameters are varied. For the tidal tests in column 1, a standard measure of non-linearity for propagating waves with $|f|<\omega$ is the steepness parameter, defined by (see e.g. Balmforth et al. (2002))
$s=kh_{0}\frac{\sqrt{N^{2}-\alpha\omega^{2}}}{\sqrt{\omega^{2}-f^{2}}}.$ (29)
For the default values in Table 1 with $|f|=5\times 10^{-5}$ and $\omega=1.4\times 10^{-4}$, setting $k$ to be the dominant wavenumber ($k=1/W$) gives $s\approx 0.6$ which is within the “subcritical” regime $s\lesssim 1$. Then, for the default steady tests in column 2 of Table 1, the Froude number is given by $Fr=0.2$ for the two-dimensional tests, and $Fr=0.3$ for the three-dimensional tests, both of which are well within the linear regime.
For efficiency, all simulations were run in hydrostatic mode, using the HydrostaticFreeSurfaceModel in Oceananigans. The unimportance of non-hydrostatic processes for the stress is reflected by the parametrisations in Section 2, whereby the hydrostatic ($\alpha=0$) and non-hydrostatic ($\alpha=1$) parametrisations are essentially the same provided the buoyancy frequency $N$ is large compared to the wave frequency. More precisely, non-hydrostatic effects are deemed unimportant for $N\gg\omega$ for tidal flows, and $N\gg kU$ for mean flows, which holds for the parameters considered here. Each three-dimensional tidal test was run for 200 hours, whereas each three-dimensional steady test was run for 100 hours. We found that 100 hours was generally long enough to reach a near-equilibrium state, with the slightly longer test time given for tidal tests as to have more full tidal cycles to perform a statistical analysis on. Since two-dimensional tests were much less computationally expensive, we decided to run them for even longer (750 hours) to improve accuracy and detect any notable effects that only appeared at later times. We timestep all simulations using a 2nd-order quasi Adams-Bashforth time stepping scheme. The free surface was substepped with the split-explicit free surface solver described by Shchepetkin and McWilliams (2005). All simulations used time-step of 7.5 minutes, and an output interval of 15 minutes.